OpenStudy (anonymous):

Need to find the formula for the area of an equilateral triangle cross section bound by the semicircles -sqrt(36-x^2) and sqrt(36-x^2)

OpenStudy (dumbcow):

 From the diagram: sin 30 = 1/2 = h/6 --> h = 3 cos 30 = sqrt3/2 = x/6 --> x = 3sqrt3 Area = x(6+h)